If the ratio 2x - y to x + y is 2/3, what is the ratio of x to y? Express your answer as a common fraction.
Answer by amarjeeth123(569)

(2x - y)/(x + y)=2/3
Simplifying we get,
6x-3y=2x+2y
6x-2x=2y+3y
4x=5y
x/y=5/4